Number of Cookies Total Utility Marginal Utility Number of Pretzels Total Utility Marginal Utility 30 30 Hide commants V2) Suppose you have a budget of $8 and cookies and pretzels cost $1 each. What is the utility maximizing combination of cooki
Answer #1

2)

Cookies MU MU/P Pretzels MU MU/P
0 0 0 0
1 25 25 1 10 10
2 15 15 2 8 8
3 10 10 3 6 6
4 5 5 4 4 4
5 0 0 5 2 2
6 -5 -5 6 0 0

The utility-maximizing combination is

4 units of each good

Total utility = 55+28 = 83

3) When the price of cookies increases to 1.5

Cookies MU MU/P Pretzels MU MU/P
0 0 0 0
1 25 16.67 1 10 10
2 15 10 2 8 8
3 10 6.67 3 6 6
4 5 3.33 4 4 4
5 0 0 5 2 2
6 -5 -3.33 6 0 0

The new utility-maximizing combination is

3 units of each good

Total utility = 50+24 = 74